OpenStudy (anonymous):

What is the projection of (2,6) onto (-1,5)?

OpenStudy (anonymous):

1.08(-1,5)

OpenStudy (amistre64):

of = b and onto = a a.b = -2+30 = 28 a.a = 2+25 = 27 28/27<-1,5> perhaps? looks to be a little over 1

OpenStudy (amistre64):

1.04 if i did it right

OpenStudy (amistre64):

1+25 aint it :)

OpenStudy (amistre64):

1.076 ... so yep